OpenStudy (anonymous):

Write an inequality for the given statement: The product of 5 and a number is at least –25.

OpenStudy (mathmate):

Hint: at least means "greater than or equal to", or \(\ge\). Example: The product of 8 and k is greater than or equal to 50. We will follow exactly the instructions: "product of 8 and k" is 8k "greater than or equal to" is \(\ge\) 50 is 50, so the whole sentence becomes 8k\(\ge\) 50.

OpenStudy (anonymous):

thx @mathmate

OpenStudy (mathmate):

You're welcome! :)